I have 4 resistors. 2 of them are 2 ohms and other 2 are 4 ohms . Now I need to arrange these resistors in a circuit so that maximum and minimum current flows through it . I need the circuit diagrams and the values of currents
Answer #1

4 2u Given o n Given a M qr VV far , maximum and minimom cunent V - when IR. v=constant. R. When R minimum Current maximum. fReq = ? I t I t 2 I tt 4 4 Reg Rey 2 + 2 + 1 + 1 4 = Reg Req= 2/3 n. cument Im = Ž 1 3 - 3 [maximum current = 3 Ampere. Ampen